In a behavioral experiment, you want to inject a 60 g crayfish with a drug and reach a drug concentration of 10-5 M in the hemolymph. The maximal volume to inject into a crayfish is 1 mL. How do you do this? How much of which concentration of the drug (in mM) do you inject? Please show work!
Answer #1

If the drug has to be injected at 10-5 M, with respect to the bodyweight of 60g, then make a stock solution of the drug, 1mM (10-3 M). Then inject 600 ul of it, so that the final concentration of the drug will be 10-5 M.

It is according to the formula :

C1 X V1=C2 X V2

Here, C1 = 10-5 M, V1 = 60 g (assume for 60ml, considering with the density of 1). C2 = 10 -3 M, and V2 = ?

So,

10-5 X 60 = 10-3 X V2

V2 = 10-2 X 60 = 0.6 mL, which is 600ul.